The height of a triangle is 12cm more than it's base. It's area is 32cm². Find the base of the triangle.
Answer by Alan3354(69443)

The height of a triangle is 12cm more than it's base. It's area is 32cm². Find the base of the triangle.
====================
After 3-4 seconds of thought,
It's 16 by 4
---
it's = it is
===================
If you wanna spend more time on it:
Area = b*h/2 = 32
b*h = 64
h = b+12
b*(b+12) = 64
b^2 + 12b - 64 = 0
Now, you find a pair of factors of 64 that differ by 12, same as above.
========================
If they were not integers, you would have to do it the "hard way."